what % of $2.50 is $1.50?
Answer by stanbon(75887)

what % of $2.50 is $1.50
-----
x/100 = 1.5/2.5
x/100 = 3/5
x = 300/5
x = 60
----------
1.50/2.50 = 60/100 = 60%
========================
Cheers,
Stan H.